# Identification of Craving Maps among Marijuana Users via the Analysis of Functional Brain Networks with High-Order Attention Graph Neural Networks

### v4

@@ Abstract
The excessive consumption of marijuana can induce substantial psychological and social consequences. In this investigation, we propose an elucidative framework termed high-order graph attention neural networks (HOGANN) for the classification of Marijuana addiction, coupled with an analysis of localized brain network communities exhibiting abnormal activities among chronic marijuana users. HOGANN integrates dynamic intrinsic functional brain networks, estimated from resting-state functional magnetic resonance imaging (rs-fMRI), using long short-term memory (LSTM) to capture temporal network dynamics. We employ a high-order attention module for information fusion and message passing among neighboring nodes, enhancing the network community analysis. Our model is validated across two distinct data cohorts, yielding substantially higher classification accuracy than benchmark algorithms. Furthermore, we discern the most pertinent subnetworks and cognitive regions affected by persistent marijuana consumption, indicating adverse effects on functional brain networks, particularly within the dorsal attention and frontoparietal networks. Intriguingly, our model demonstrates superior performance in cohorts exhibiting prolonged dependence, implying that prolonged marijuana usage induces more pronounced alterations in brain networks. The model proficiently identifies craving brain maps, thereby delineating critical brain regions for analysis.
Keywords Marijuana  $\cdot$ fMRI  $\cdot$ Graph neural network (GNN)  $\cdot$ Multigraph classification  $\cdot$ Addiction prediction
@@ 1 Introduction
The issue of substance use disorder (SUD) has become a pressing concern in the United States, with the legalization of non-medical marijuana usage. Marijuana-related disorder accounts for a substantial proportion of individuals seeking treatment for drug use disorders due to the high global prevalence of marijuana use [1]. Heavy marijuana consumption has been found to impact brain function and cognition  [2, 3, 4]. In the traditional view, addiction is associated with brain abnormalities in the nucleus accumbens, prefrontal cortex, and amygdala, extensively investigated through neuroimaging techniques [5, 6, 7]. The regional structural or functional changes among patients with SUD may influence addictive behavior by altering reward processing, decision-making, and emotional regulation [8]. However, current research indicates that addiction is not merely a localized brain abnormality but a disruption at the network level. Interacting neural assemblies or brain regions constitute brain networks, and disruptions in connectivity between different regions may crucially impact addictive behavior. Addiction is thus viewed as a network-wide phenomenon with altered functional connectivity patterns, as indicated by network analysis based on fMRI [9, 10, 11, 12]. Recent studies have shown that leveraging fMRI-derived brain network analysis can offer new insights into the relationships between underlying brain connectivity alterations and the manifested characteristics of addictive behavior [13, 14, 15]. These characteristics include the age of onset of marijuana use, environmental influences, symptoms of addiction [14], and predictions of marijuana withdrawal and personalized treatment [13].
@@
Constructing brain connectomes using functional or structural neuroimaging modalities has become one of the most pervasive frameworks for brain imaging analysis [20]. In recent years, graph neural networks (GNNs) have been employed to analyze brain networks for neurological disorder classification and subtyping [20]. These functional brain networks are constructed from correlation maps of fMRI, where each vertex represents a region of interest (ROI), and edges represent the connection strength between different ROIs  [21, 22, 23, 24]. Researchers have utilized a multimodal approach of diffusion tensor imaging (DTI) and fMRI to diagnose mental network disorders [22], and have also built functional brain networks to classify autism spectrum disorder (ASD)  [25, 26, 27]. Moreover, Schizophrenia can be identified with salient regions based on its topological structure using SVM and Graph Convolutional Network (GCN), as demonstrated by six rs-fMRI datasets. [28].
However, fewer studies have conducted functional brain network analysis based on GNN for long-term marijuana users to (1) improve the classification accuracy and (2) delineate the related craving maps or explain alterations of brain networks. To bridge the research gap, in this study, we propose an interpretable high-order attention graph attention neural networks (HOGANN) model for analyzing the abnormal brain activities demonstrated in long-term marijuana users compared to healthy controls (HC). The main contributions of this study can be summarized as follows: (1) proposing to utilize graph deep learning methods to classify long-term marijuana users, (2) leveraging the sequential multigraphs in the fMRI time series of marijuana users, (3) integrating the functionality of high-order nodes for information fusion and message passing into our model, and (4) providing improved clustering analysis for identifying craving maps among long-term marijuana users. Furthermore, we compared multiple benchmark methods and evaluated the effectiveness of the proposed HOGANN. By training the model and reconstructing weighted brain networks, we found that the HOGANN is more effective at analyzing differences between long-term marijuana users and non-users as characterized by functional brain connectivity networks.
@@ 2.1 Problem Definition
We define the classification of SUD for each subject as a supervised multigraph classification problem, where each instance graph is constructed from a sliding window of the resting state fMRI (rs-fMRI) time series. Specifically, we denote the multivariate time-series data for all individuals as $\{\textbf{X}^{(i)}\}_{i=1}^{N}$ with each individual’s data being $\textbf{X}^{(i)}\in\mathbb{R}^{v\times T}$, where $v$ represents the number of ROIs, and $T$ is the time series length. We segment these non-overlapping time series into multiple segments, each with a length of $T^{\prime}$. This results in a total of $\left|K\right|$ segments for each subject. Subsequently, we construct the multigraph $\{\textbf{G}^{(i)}(E,V;t)\}_{t=1}^{K}$ = $\{\textbf{A}^{(i)}(t)\}_{t=1}^{K}$ from the segmented subsequences, where the functional connectivity is based on the k-nearest neighbor (k-NN) graph of each ROI for each subject, where $E$ and $V$ denotes edges and vertex (region) set, respectively. Each adjacency matrix $\textbf{A}^{(i)}(t)\in\mathbb{R}^{v\times v}$ corresponds to the $k$th segment in the $t$th time window denoted $X^{(i)}(t),(i\in 1,\cdots,N,t\in 1,\cdots,K)$, and corresponding class labels $Y^{(i)}$ for the multigraph classification task.
@@ 2.2 Model Architecture
The overall framework of HOGANN, as shown in Fig. 1encompasses a multigraph classification architecture with two sub-models: (1) the high-order graph attention neighbor mixing time series model, which captures features from multigraphs of the same subject, and (2) the sequential dynamic graph learning model using LSTM to learn the temporal interdependence of graphs. The high-order graph attention neighbor mixing module employs a higher-order attention mechanism with a neighborhood mixing model to learn the relationship between ROIs after segmenting dynamic time series in the same subject. Meanwhile, the proposed LSTM sub-model can better extract local features from the sequential graphs to improve distinguishability between marijuana subjects and normal controls. The overall framework enables concatenating features from the latent spaces of the two sub-models leading to improved model performance.
Figure 1: The overall framework of the HOGANN utilizes two sub-models to perform fusion learning on fMRI time series. The high-order graph attention neighbor mixing model combines hopping to conduct high-order attention for message passing. Sequential dynamic graph learning uses LSTM to learn the temporal sequence graph and discern differences between instances.
@@ 2.2.1 High-order graph attention neighbor mixing
To learn dynamic topological brain networks from multigraphs, we utilize one-hop connections for local graph node message passing with a MixHop layer [29]. The formulated equation of the classical $l$-th graph hidden convolutional layer can be defined as follows:
@@
where normalized adjacency matrix is given by $\hat{\textbf{A}}=\textbf{D}^{-\frac{1}{2}}(\textbf{A}+\textbf{I}_{N})\textbf{D}^{-\frac{1}{2}}$, and $\textbf{A}+\textbf{I}_{N}$ denotes the self-loop connection; $D_{ii}=\sum_{i}A_{ii}$ presents the degree matrix of A, W is a learnable weight matrix in the non-linear activation function $\sigma(\cdot)$. By incorporating the neighborhoods of high-order nodes, the MixHop layer facilitates higher-order latent feature learning enables learning from their neighbors by incorporating the $j$-th power of the self-adjacency matrix $\hat{\textbf{A}}^{j}$, and the two-hop delta operator learns for various neighborhood distances by subtracting different node features. The MixHop graph convolutional layer can be defined as
$\textbf{M}^{(l+1)}=\mathop{\Big{\|}}\limits_{j\in P}\sigma(\hat{\textbf{A}}^{j}\textbf{M}^{(i)}\textbf{W}_{j}^{(i)}),$ (2)
@@
Attention with MixHop fusion: To effectively integrate temporal graph features and the strength of connectivity paths between nodes through the attention mechanism, we utilized both the MixHop architecture employing high-order message passing and the GAT module for fusing multigraph node features, as depicted in Fig. 2. The GAT layer, with graph attention mechanisms, aggregate neighborhood node features with weighted attention score, while the MixHop layer enhances the connectivity path from source node to high neighbor node of functional connectivity [30]. (Note: first-degree to $k$-degree proximity). We denote the node embedding $\textbf{H}^{0}=\{\vec{h}^{0}_{1},\vec{h}^{0}_{2},...,\vec{h}^{0}_{n}\}$ from the first graph attention layer and learn the neighbors $\mathcal{N}_{i}$ of ROI nodes $v_{j}$. We can update the parameterized feature vectors $h_{i}$ and the learnable weights $\textbf{W}^{0}\in\mathbb{R}^{n\times d}$ using the LeakyReLU function to further calculate the importance attention score as Eq. (3)
@@
$\alpha_{ij}=\frac{exp\left(LeakyReLU(\vec{a}^{T}_{ij}\left[\textbf{W}\vec{h}_{i}\mathbin{\|}\textbf{W}\vec{h}_{j}\right]\right)}{\sum_{k\in\mathcal{N}_{i}}exp\left(LeakyReLU(\vec{a}^{T}_{ij}\left[\textbf{W}\vec{h}_{i}\mathbin{\|}\textbf{W}\vec{h}_{k}\right]\right)},$ (3)
@@
The attention mechanism generates an importance attention score $\alpha_{ij}\in\mathbb{R}$ for each edge between ROIs, enabling the neighboring vertex to better focus on the corresponding node at the edge. Subsequently, we can calculate the features representation $\hat{h}$ of the corresponding nodes through the attention coefficients after a non-linear transformation, as follows:
$\hat{h}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\textbf{W}\vec{h}_{j}\right),$ (4)
@@
where $\sigma(\cdot)$ denotes a nonlinear transformation function. In the second layer of the sub-model, we enhance the attention mechanism and the capability for high-order feature fusion within the overall network, as illustrated in Fig. 2. We concatenate embedding vectors $m$ and $\hat{h}$ from the MixHop and GAT layers as the input to the second-layer graph attention layer. However, traditional attention weights for each edge are primarily effective for learning first-order node connectivity, whereas attention MixHop fusion excels at learning second-order neighbor proximity pathways. The concatenated embedding is utilized as the aggregation input for the second-layer $\textbf{GAT}_{l+1}$, as follows:
$\textbf{H}^{a}=\textbf{GAT}_{l+1}(\phi(\left[\textbf{M}\oplus\hat{\textbf{H}}\right])),$ (5)
@@
where M and $\textbf{H}^{a}=\left[\hat{h}_{1}^{a},\hat{h}_{2}^{a},...,\hat{h}_{k}^{a}\right]$ are output features vectors, $\oplus$ present concatenation with corresponding output channel dimensions, $\phi(\cdot)$ denotes graph global mean pooling. The $\frac{1}{\left|\mathcal{V}\right|}\sum_{i\in\mathcal{V}}\hat{h}^{a(l)}_{i}$ indicates averaging node features across the node dimension $\left|\mathcal{V}\right|$.
2.2.2 Sequential graph learning
@@
Figure 2: The proposed first sub-model contains MixHop and a graph attention layer, wherein the second layer of message passing, concatenating different powers of $\hat{A}$, can enhance the depth of attention in the neighborhood node connectivity in the pathway.
@@
The fMRI signals of the brain’s ROI exhibit variations in dynamic network representation over time $T$ for each instance. The second sub-model was designed to extract dynamic features of functional graphs with cross-sequential dependencies using a batch of $b$ multigraphs ${\{\textbf{A}^{(i)}}\}_{i=1}^{N}$, where $b\subseteq|K|\times|N|$. LSTM networks have found extensive application in handling sequential data, particularly in scenarios where extracting interrelated sequential time series at various time points is necessary. Starting from time step t, we can formulate the sequence learning representation, which comprises the input gate $t$, we can formulate the sequence learning representation, which comprises the input gate $\textbf{I}_{t}$, forget gate $\textbf{F}_{t}$, output gate $\textbf{O}_{t}$, cell gate $\textbf{C}_{t}$, and output gate $\textbf{H}^{g}$, as depicted:
Figure 3: The proposed second sub-model, LSTM, aims to train the graph of each batch instance and learn representations of spatial-temporal dependencies for different subjects.
@@
$\tilde{\textbf{C}}_{t}$ $=tanh(\textbf{W}\cdot\textbf{GATconv}(\textbf{A}^{(i)}(t))+\textbf{W}\cdot\textbf{GATconv}(H_{t-1}))$ (9)
$\textbf{C}_{t}$ $=tanh(\textbf{I}_{t}\cdot\tilde{\textbf{C}}_{t}+\textbf{F}_{t}\cdot\textbf{C}_{t-1})$ (10)
H $=\textbf{O}_{t}\cdot tanh(\textbf{C}_{t})$ (11)
@@
where $\textbf{W}{i}$, $\textbf{W}_{f}$, $\textbf{W}{o}$ and $\textbf{W}_{c}$ represent learnable weights at different gates and bias vectors $b_{i}$, $b_{f}$, and $b_{o}$. By taking the output from the LSTM layer, we obtain the final instance graph representation:
$h^{\prime}_{1},h^{\prime}_{2},...,h^{\prime}_{k}=\textbf{LSTM}(h_{1},h_{2},...,h_{k}),$ (12)
@@
where $h^{\prime}_{1},h^{\prime}_{2},...,h^{\prime}_{k}$ are latent vectors in $t$-th time step from LSTM modeling as illustrated in Fig. 3. Then, we can combine temporal graph features and use graph mean pooling $\phi(\cdot)$ to generate instance representation $\textbf{H}^{\prime}=\phi(\left[h_{1}^{\prime}\lVert h_{2}^{\prime}\lVert h_{3}^{\prime}...\lVert h_{k}^{\prime}\right])$. To better aggregate two sub-model embeddings, we combine two representations of $\textbf{H}^{a}$ and $\textbf{H}^{\prime}$ for the final embedding output as shown:
$\textbf{H}^{c}_{i}=\textbf{H}^{a}+\textbf{H}^{\prime}.$ (13)
@@
where $\textbf{H}^{c}_{i}\in\mathbb{R}^{B\times C}$ represents the final output, and C denotes the final prediction class. This prediction class is a combination of outputs from the two sub-models and is utilized to integrate information contributed by both perspectives provided by the sub-models.
2.2.3 Brain network reconstruction and community detection

### v5

# Spatial Craving Patterns in Marijuana Users: Insights from fMRI Brain Connectivity Analysis with High-Order Graph Attention Neural Networks

@@ Abstract
The excessive consumption of marijuana can induce substantial psychological and social consequences. In this investigation, we propose an elucidative framework termed high-order graph attention neural networks (HOGANN) for the classification of Marijuana addiction, coupled with an analysis of localized brain network communities exhibiting abnormal activities among chronic marijuana users. HOGANN integrates dynamic intrinsic functional brain networks, estimated from functional magnetic resonance imaging (fMRI), using graph attention-based long short-term memory (GAT-LSTM) to capture temporal network dynamics. We employ a high-order attention module for information fusion and message passing among neighboring nodes, enhancing the network community analysis. Our model is validated across two distinct data cohorts, yielding substantially higher classification accuracy than benchmark algorithms. Furthermore, we discern the most pertinent subnetworks and cognitive regions affected by persistent marijuana consumption, indicating adverse effects on functional brain networks, particularly within the dorsal attention and frontoparietal networks. Intriguingly, our model demonstrates superior performance in cohorts exhibiting prolonged dependence, implying that prolonged marijuana usage induces more pronounced alterations in brain networks. The model proficiently identifies craving brain maps, thereby delineating critical brain regions for analysis.
Keywords Marijuana  $\cdot$ fMRI  $\cdot$ Graph neural network (GNN)  $\cdot$ Multigraph classification  $\cdot$ Addiction prediction  $\cdot$ Brain connectivity analysis
@@ 1 Introduction
The issue of substance use disorder (SUD) has become a pressing concern in the United States, with the legalization of non-medical marijuana usage. Marijuana-related disorder accounts for a substantial proportion of individuals seeking treatment for drug use disorders due to the high global prevalence of marijuana use [1]. Heavy marijuana consumption has been found to impact brain function and cognition  [2, 3, 4]. In the traditional view, addiction is associated with brain abnormalities in the nucleus accumbens, prefrontal cortex, and amygdala, extensively investigated through neuroimaging techniques [5, 6, 7, 8]. The regional structural or functional changes among patients with SUD may influence addictive behavior by altering reward processing, decision-making, and emotional regulation [9]. However, current research indicates that addiction is not merely a localized brain abnormality but a disruption at the network level. Interacting neural assemblies or brain regions constitute brain networks, and disruptions in connectivity between different regions may crucially impact addictive behavior. Addiction is thus viewed as a network-wide phenomenon with altered functional connectivity patterns, as indicated by network analysis based on fMRI [10, 11, 12, 13]. Recent studies have shown that leveraging fMRI-derived brain network analysis can offer new insights into the relationships between underlying brain connectivity alterations and the manifested characteristics of addictive behavior [14, 15, 16]. These characteristics include the age of onset of marijuana use, environmental influences, symptoms of addiction [14], and predictions of marijuana withdrawal and personalized treatment [14].
@@
Constructing brain connectomes using functional or structural neuroimaging modalities has become one of the most pervasive frameworks for brain imaging analysis [21]. In recent years, graph neural networks (GNNs) have been employed to analyze brain networks for neurological disorder classification and subtyping [21]. These functional brain networks are constructed from correlation maps of fMRI, where each vertex represents a region of interest (ROI), and edges represent the connection strength between different ROIs  [22, 23, 24, 25]. Researchers have utilized a multimodal approach of diffusion tensor imaging (DTI) and fMRI to diagnose mental network disorders [23], and have also built functional brain networks to classify autism spectrum disorder (ASD)  [26, 27, 28]. Moreover, Schizophrenia can be identified with salient regions based on its topological structure using SVM and Graph Convolutional Network (GCN) [29], as demonstrated by six fMRI datasets [30].
However, fewer studies have conducted functional brain network analysis based on GNN for long-term marijuana users to (1) improve the classification accuracy and (2) delineate the related craving maps or explain alterations of brain networks.
To bridge the research gap, in this study, we propose an interpretable high-order graph attention neural networks (HOGANN) model for analyzing the abnormal brain activities demonstrated in long-term marijuana (LM) users compared to healthy controls (HC). The main contributions of this study can be summarized as follows: (1) proposing to utilize graph deep learning methods to classify LM users, (2) leveraging the sequential multigraphs in the fMRI time series of marijuana users, (3) integrating the functionality of high-order attention for node relation awareness and message passing into our model, and, (4) we employed analysis across two data sets, successfully identifying significant subnetworks associated with craving maps among LM users, which provided valuable insights into the underlying mechanisms.
Figure 1: The overall framework of the HOGANN utilizes two sub-models to perform fusion learning on fMRI time series. The high-order graph attention neighbor mixing model combines hopping to conduct high-order attention for message passing. Sequential graph attention learning uses GAT-LSTM to learn the temporal sequence graph and discern differences between instances.
@@ 2.1 Problem Definition
We define the classification of SUD for each subject as a supervised multigraph classification problem, where each instance graph is constructed from a sliding window of the fMRI time series. Specifically, we denote the multivariate time-series data for all individuals as $\{\textbf{X}^{(i)}\}_{i=1}^{N}$ with each individual’s data being $\textbf{X}^{(i)}\in\mathbb{R}^{v\times T}$, where $v$ represents the number of ROIs, and $T$ is the time series length. We segment these non-overlapping time series into multiple segments, each with a length of $T^{\prime}$. This results in a total of $\left|K\right|$ segments for each subject. Subsequently, we construct the multigraph $\{\textbf{G}^{(i)}(E,V;t)\}_{t=1}^{K}$ = $\{\textbf{A}^{(i)}(t)\}_{t=1}^{K}$ from the segmented subsequences, where the functional connectivity is based on the k-nearest neighbor (k-NN) graph of each ROI for each subject, where $E$ and $V$ denote edges and vertex (region) set, respectively. Each adjacency matrix $\textbf{A}^{(i)}(t)\in\mathbb{R}^{v\times v}$ corresponds to the $t$-th segment in the $i$-th subject’s time series, denoted $\textbf{X}^{(i)}(t),(i\in 1,\cdots,N,t\in 1,\cdots,K)$, and corresponding class labels $Y^{(i)}$ for the multigraph classification task.
@@ 2.2 Model Architecture
The overall framework of HOGANN, as shown in Fig. 1 encompasses a multigraph classification architecture with two sub-models: (1) the high-order graph attention neighbor mixing dynamic model, which captures features from multigraphs representations of subject fMRI time series, and (2) the sequential graph attention learning model utilizing graph attention-based long short-term memory (GAT-LSTM) to capture the temporal interdependencies among graphs. The advanced graph attention neighbor mixing module integrates graph neighborhood node message passing to enhance learning of inter-regional relationships of ROIs among individual subjects. Meanwhile, the proposed GAT-LSTM sub-model enhances the extraction of structural features from sequential graphs of interconnected nodes, thereby improving the ability to differentiate between marijuana users and healthy controls. The overall framework enables concatenating features from the latent spaces of the two sub-models leading to improved model performance.
2.2.1 Two-hop Delta Operator
We can define injective mapping Delta Operator $f$ to learn distance between two-hop node features from first-degree neighbors to second-degree neighbors, as follow:
$f\left(\sigma\left(\hat{A}X\right)-\sigma\left(\hat{A}^{2}X\right)\right),$ (1)
where operation can be given any $\hat{A}$, input features $X$ and activate function $\sigma$. The operator allows the model to represent feature differences among neighbors, enabling it to learn the boundary conditions of the graph.
@@ 2.2.2 High-order graph attention neighbor mixing
To learn dynamic topological brain networks from multigraphs, we can directly replace the traditional graph convolution layers from the transformation of the two-hop Delta Operation based on the Eq. (1). We utilize one-hop connections for local graph node message passing with a MixHop layer [31]. The formulated equation of the classical $l$-th graph hidden convolutional layer can be defined as follows:
@@
where normalized adjacency matrix is given by $\hat{\textbf{A}}=\textbf{D}^{-\frac{1}{2}}(\textbf{A}+\textbf{I}_{N})\textbf{D}^{-\frac{1}{2}}$, and $\textbf{A}+\textbf{I}_{N}$ denotes the self-loop connection; $D_{ii}=\sum_{j}A_{ij}$ presents the diagonal degree matrix of A, $\textbf{W}^{(l)}$ is a learnable weight matrix in the non-linear activation function $\sigma(\cdot)$. By incorporating the neighborhoods of high-order nodes, the MixHop layer facilitates higher-order latent feature learning enables learning from their neighbors by incorporating the $j$-th power of the self-adjacency matrix $\hat{\textbf{A}}^{j}$, and the two-hop delta operator learns for various neighborhood distances by subtracting different node features. The MixHop graph convolutional layer can be defined as
$\textbf{M}^{(l+1)}=\mathop{\Big{\|}}\limits_{j\in P}\sigma(\hat{\textbf{A}}^{j}\textbf{M}^{(l)}\textbf{W}_{j}^{(l)}),$ (3)
@@
Attention with MixHop fusion: To effectively integrate temporal graph features and the strength of connectivity paths between nodes through the attention mechanism, we utilized both the MixHop architecture employing high-order message passing and the graph attention network (GAT) [32] module for fusing multigraph node features, as depicted in Fig. 2. The GAT layer, with graph attention mechanisms, aggregate neighborhood node features with weighted attention score, while the MixHop layer enhances the connectivity path from source node to high neighbor node of functional connectivity [33]. (Note: first-degree to $k$-degree proximity). We denote the initial node embedding $\textbf{H}^{0}=[\vec{h}^{0}_{1},h^{0}_{2},...,\vec{h}^{0}_{n}]$ from the first graph attention layer and learn the neighbors $\mathcal{N}_{i}$ of ROI nodes $v_{j}$. We can update the parameterized feature vectors $\vec{h}_{i}$ and the learnable weights $\textbf{W}\in\mathbb{R}^{v\times d}$ using the LeakyReLU function to further calculate the importance attention score with parameterized weight vector $\vec{\textbf{a}}$ across neighboring node $j$ as:
@@
$\alpha_{ij}=\frac{\mathrm{exp}\left(\mathrm{LeakyReLU}(\vec{\textbf{a}}^{T}\left[\textbf{W}\vec{h}_{i}\mathbin{\|}\textbf{W}\vec{h}_{j}\right]\right)}{\sum_{k\in\mathcal{N}_{i}}\mathrm{exp}\left(\mathrm{LeakyReLU}(\vec{\textbf{a}}^{T}\left[\textbf{W}\vec{h}_{i}\mathbin{\|}\textbf{W}\vec{h}_{k}\right]\right)},$ (4)
@@
The attention mechanism generates an importance attention score $\alpha_{ij}$ for each edge between ROIs, enabling the neighboring vertex to better focus on the corresponding node at the edge. Subsequently, we can calculate the features representation $\hat{\textbf{H}}=[\hat{h}_{1},\hat{h}_{2},...,\hat{h}_{v}]$ of the corresponding $v$ nodes through the attention coefficients after a non-linear transformation, as follows:
$\hat{h}_{i}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\textbf{W}\vec{h}_{j}\right),$ (5)
@@
where $\sigma(\cdot)$ denotes a nonlinear transformation function. In the second layer of the sub-model, we enhance the attention mechanism and the capability for high-order feature fusion within the overall network, as illustrated in Fig. 2. Furthermore, we concatenate the embeddings M and $\hat{\textbf{H}}$ from the MixHop and GAT layers, respectively, as the input to the second-layer graph attention layer. However, traditional attention weights for each edge are primarily effective for learning first-order node connectivity, whereas attention MixHop fusion excels at learning second-order neighbor proximity pathways. The concatenated embedding is utilized as the aggregation input for the second-layer $\textbf{GAT}^{(l+1)}$, as follows:
$\textbf{H}^{a}=\textbf{GAT}^{(l+1)}(\left[\textbf{M}\oplus\hat{\textbf{H}}\right]),$ (6)
@@
where $\textbf{H}^{a}=\left[\hat{h}_{1}^{a},\hat{h}_{2}^{a},...,\hat{h}_{k}^{a}\right]$ are output aggregate features vectors, $\oplus$ presents concatenation with corresponding output channel dimensions.
2.2.3 Attention-driven sequential multigraph learning
@@
Figure 2: The proposed first sub-model contains MixHop and a graph attention layer, wherein the second layer of message passing, concatenating different powers of $\hat{A}^{j}$, can enhance the depth of attention in the neighborhood node connectivity in the pathway.
@@
The fMRI signals of the brain’s ROIs exhibit variations in dynamic network representation over time $T$ for each instance. The second sub-model was designed to extract dynamic features of functional graphs with cross-sequential dependencies using sequential multigraphs $\{\textbf{A}^{(i)}(t)\}_{t=1}^{K}$. Traditional LSTM does not consider the graph structure dependencies in longer time sequences. To address this limitation, we propose a modified LSTM architecture that incorporates graph attention mechanisms to capture spatial correlations in node embeddings for multigraph sequences. This attention mechanism effectively enhances the model’s ability to learn intrinsic sequential graph structures, particularly in scenarios where extracting interrelated sequential time series at various time points is necessary. Starting from time step $t$, we can formulate the sequence learning from temporal graph representations using the following components: input gate $\textbf{I}_{t}$, forget gate $\textbf{F}_{t}$, output gate $\textbf{O}_{t}$, modulated input $\tilde{\textbf{C}}_{t}$, cell gate $\textbf{C}_{t}$, and output gate $\textbf{H}_{t}$, as depicted in the following equations:
@@
$\tilde{\textbf{C}}_{t}$ $=\tanh(\textbf{W}_{c}\cdot\textbf{GATconv}(\textbf{A}^{(i)}(t))+\textbf{W}_{c}\cdot\textbf{GATconv}(H_{t-1}))$ (10)
$\textbf{C}_{t}$ $=\tanh(\textbf{I}_{t}\cdot\tilde{\textbf{C}}_{t}+\textbf{F}_{t}\cdot\textbf{C}_{t-1})$ (11)
$\textbf{H}_{t}$ $=\textbf{O}_{t}\cdot\tanh(\textbf{C}_{t})$ (12)
where $\textbf{W}_{i}$, $\textbf{W}_{f}$, $\textbf{W}{o}$ and $\textbf{W}_{c}$ represent learnable weights at different gates and bias vectors $b_{i}$, $b_{f}$, and $b_{o}$. Notably, we introduce the graph attention convolution layer GATconv()($\cdot$)( ⋅ ) to encode sequential graphs $\textbf{A}^{(i)}(t)$ in each LSTM gate. We can define $l$-th GATconv()($\cdot$)( ⋅ ) layer as:
$\textbf{H}_{t}^{(l+1)}=\textbf{GATconv}(\textbf{A}^{(i)}(t))=\sigma(\textbf{A}^{(i)}(t)\textbf{H}^{(l)}_{t}\textbf{W}^{(l)}_{t}),$ (13)
@@
where $\textbf{H}^{(l+1)}_{t}$ is the output features matrix calculated after the $l$-th hidden layer of graph attention with trainable weights $\textbf{W}^{(l)}_{t}$. By taking the output from the GAT-LSTM layer, we obtain the final graph representation for each instance:
$h^{\prime}_{1},h^{\prime}_{2},...,h^{\prime}_{k}=\textbf{GAT-LSTM}(h_{1},h_{2},...,h_{k}),$ (14)
@@
where $h^{\prime}_{1},h^{\prime}_{2},...,h^{\prime}_{k}$ are latent vectors in $t$-th time step from graph attention LSTM modeling as illustrated in Fig. 1. Then, we can combine temporal graph features to generate instance representation $\textbf{H}^{\prime}=\left[h_{1}^{\prime},h_{2}^{\prime},h_{3}^{\prime},...,h_{k}^{\prime}\right]$. To better aggregate two sub-model embeddings, we first combine the representations $\textbf{H}^{a}$ and $\textbf{H}^{\prime}$, as follows:
$\textbf{H}^{c}=\frac{1}{2}(\textbf{H}^{a}+\textbf{H}^{\prime}).$ (15)
@@
where $\textbf{H}^{c}$ represents final fused node embedding, which is a combination of outputs from the two sub-models and is utilized to integrate information contributed by both perspectives provided by these sub-models.
2.2.4 Brain network reconstruction and weighted functional connectivity
Functional connectivity assesses brain connections between various regions. To improve brain network connectivity among different ROI/nodes with spatial relationships, we calculate node $i$ features representation using a mixhop attention mechanism by aggregating fused node features from Eq. 15. Subsequently, we perform graph mean pooling for $i$th node embeddings, which can be expressed as:
$L^{c}_{i}=\frac{1}{\left|\mathcal{V}\right|}\sum_{j\in\mathcal{V}}H^{c}_{j},$ (16)